OpenStudy (anonymous):

Factor completely: ax + 2ay + bx + 2by

OpenStudy (anonymous):

\[(a+b)(x+2y)\]

Directrix (directrix):

ax + 2ay + bx + 2by = a ( x + 2y) + b ( x + 2y ) = (x + 2y) (a + b) Factoring by Grouping